# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 6.3.1 Color image denoising without NSS
Image Algorithm SRRC[39] WNNM[14] DWT[5] LRQA-1[10] LRQA-2[10] SLRQA-1 SLRQA-2 SLRQA-3
$\tau=10$ Image1 29.23/0.70 28.06/0.73 27.67/0.61 28.48/0.61 29.21/0.65 31.55/0.76 31.45/0.78 32.86/0.81
Image2 28.56/0.73 28.01/0.70 27.98/0.71 28.44/0.72 28.42/0.72 30.36/0.80 30.19/0.80 31.20/0.82
Image3 28.03/0.60 27.81/0.61 28.21/0.61 28.49/0.64 28.56/0.65 31.33/0.77 31.79/0.78 32.03/0.84
Image4 28.12/0.60 27.90/0.65 28.29/0.66 28.53/0.64 28.55/0.68 30.09/0.74 31.45/0.75 30.36/0.75
Image5 28.82/0.66 28.72/0.64 28.67/0.64 28.85/0.65 28.71/0.67 32.13/0.85 31.18/0.80 32.50/0.87
Image6 28.32/0.60 27.89/0.56 28.02/0.61 28.72/0.65 28.66/0.65 30.75/0.76 30.92/0.80 30.77/0.76
Image7 28.45/0.62 27.32/0.58 28.02/0.59 28.65/0.60 29.04/0.94 30.95/0.73 30.82/0.75 31.10/0.80
Image8 28.01/0.60 28.11/0.59 28.46/0.58 28.56/0.63 28.24/ 0.63 31.23/0.76 30.75/0.74 31.92/0.76
Image9 28.02/0.74 27.98/0.73 27.62/0.71 28.82/0.76 29.10/0.76 30.14/0.70 29.66/0.78 30.80/0.79
Image10 30.52/0.74 29.72/0.72 29.62/0.70 30.90/0.76 31.10/0.79 31.14/0.78 31.26/0.80 31.32/0.81
$\tau=30$ Image1 20.32/0.30 20.41/0.25 20.36/0.28 20.49/0.27 20.80/0.30 25.47/0.66 24.30/0.61 26.40/0.67
Image2 20.01/0.32 20.73/0.32 20.09/0.33 20.44/0.37 20.40/0.37 23.90/0.61 22.77/0.64 24.02/0.64
Image3 00.12/0.21 20.24/0.22 20.55/0.25 20.50/0.27 21.10/0.31 26.15/0.65 25.21/0.64 26.22/0.68
Image4 20.22/0.29 20.10/0.26 20.01/0.29 20.68/0.29 20.20/0.31 23.81/0.59 22.74/0.57 24.21/0.64
Image5 21.02/0.32 20.58/0.33 20.65/0.37 21.23/0.30 21.30/0.33 26.60/0.68 25.32/0.58 26.34/0.68
Image6 20.87/0.33 21.01/0.37 20.37/0.37 20.76/0.36 20.61/0.35 23.95/0.66 21.72/0.64 22.50/0.63
Image7 20.33/0.29 20.71/0.29 20.94/0.22 20.93/0.26 20.21/0.21 25.12/0.63 24.39/0.68 26.42/0.68
Image8 20.56/0.22 20.42/0.24 21.08/0.28 20.77/0.29 20.90/0.29 25.14/0.64 24.06/0.62 26.12/0.67
Image9 24.82/0.58 24.22/0.58 23.62/0.55 24.90/0.58 25.10/0.61 25.30/0.64 25.45/0.68 25.80/0.68
Image10 24.32/0.64 23.82/0.62 23.62/0.61 24.34/0.66 24.90/0.68 24.97/0.66 25.02/0.70 25.68/0.70
$\tau=50$ Image1 17.92/0.17 17.01/0.14 18.01/0.18 17.66/0.17 17.30/0.18 23.33/0.50 23.50/0.51 23.90/0.51
Image2 16.02/0.21 17.03/0.20 17.94/0.25 17.53/0.25 17.10/0.26 22.04/0.48 22.77/0.52 22.06/0.45
Image3 17.75/0.15 17.40/0.21 17.24/0.19 17.56/0.17 17.90/0.18 24.21/0.51 24.91/0.58 24.99/0.60
Image4 17.08/0.14 16.60/0.16 16.87/0.17 17.82/0.19 17.20/0.14 22.23/0.46 22.34/0.47 23.51/0.49
Image5 18.04/0.17 18.08/0.17 18.31/0.18 18.39/0.19 18.30/0.17 21.84/0.52 21.62/0.67 22.72/0.65
Image6 17.53/0.25 18.01/0.26 17.11/0.22 17.75/0.26 18.61/0.29 21.24/0.50 21.02/0.51 22.00/0.53
Image7 17.60/0.17 18.21/0.18 17.11/0.13 18.08/0.17 18.71/0.17 23.57/0.48 23.09/0.45 23.92/0.53
Image8 17.02/0.14 17.42/0.18 17.62/0.19 17.90/0.18 18.10/0.29 23.14/0.48 23.26/0.52 24.32/0.56
Image9 22.02/0.51 21..92/0.49 21.62/0.46 22.34/0.52 22.67/0.55 22.78/0.56 23.02/0.62 23.04/0.63
Image10 21.82/0.50 21.42/0.48 21.32/0.45 22.32/0.51 22.10/0.51 22.74/0.54 22.60/0.52 23.10/0.58
@@
(c) SRRC
(d) WNNM
(e) DWT
(f) LRQA-1
(g) LRQA-2
(h) SLRQA-1
(i) SLRQA-2
(j) SLRQA-3
Figure 4: Color image denoising results of different methods on “Image5” with $\tau=10$.
@@ 6.3.2 Color image denoising with NSS
For color image denoising problems with NSS, in addition to the methods compared in Section 6.3.1, we also compare SLRQA with CBM3D [11] [^4] The implementation is available from http://www.cs.tut.fi/~foi/GCF-BM3D.. Instead of constructing a denoising model, CBM3D employs block matching and sparse 3D collaborative filtering techniques to denoise the image, a process that cannot be achieved without using NSS. Hence we do not conduct a comparison with the CBM3D algorithm in the experiments of Section 6.3.1.
@@
The PSNR and SSIM values of each image between the competing methods under different noise levels are reported in Table 5. Most proposed SLRQA methods achieve higher PSNR and SSIM values than other methods. Moreover, SLRQA-3 achieves the best results in most cases. The visual comparisons of Image5 with $\tau=30$ are shown in Figure 6 as a typical example. It is shown in Figure 6 that the images recovered by CBM3D and WNNM methods are still noisy. In addition, the images recovered by DWT, LRQA-1, and LRQA-2 are overly smooth, resulting in the loss of the features of the original images. In contrast, the image recovered by SLRQA-3 can not only remove the noise but also keep the features of the original images.
@@
Table 5: PSNR/SSIM results on color image denoising problem of different methods with NSS.
@@
Image Algorithm CBM3D[11] SRRC[39] WNNM[14] DWT[5] LRQA-1[10] LRQA-2[10] SLRQA-1 SLRQA-2 SLRQA-3
$\tau=10$ Image1 35.71/0.92 35.67/0.92 34.85/0.91 34.62/0.91 35.19/0.92 35.34/0.92 35.69/0.92 35.57/0.92 35.86/0.93
Image2 32.22/0.86 32.02/0.85 31.44/0.86 30.89/0.85 31.94/0.87 31.88/0.87 32.20/0.87 32.19/0.87 32.40/0.87
Image3 34.67/0.88 34.72/0.88 34.37/0.88 34.22/0.88 34.63/0.88 34.72/0.88 34.75/0.88 34.79/0.88 35.03/0.88
Image4 34.36/0.91 34.12/0.91 33.95/0.91 33.89/0.91 34.01/0.91 34.05/0.91 34.43/0.91 34.45/0.91 34.36/0.91
Image5 37.08/0.95 36.82/0.94 35.94/0.95 35.87/0.94 36.48/0.95 36.71/0.95 37.10/0.95 37.18/0.95 37.50/0.95
Image6 37.02/0.97 36.32/0.96 36.04/0.96 36.02/0.96 36.65/0.97 36.66/0.97 37.39/0.97 37.22/0.97 37.27/0.97
Image7 34.25/0.91 34.01/0.89 33.80/0.90 33.92/0.90 34.25/0.90 34.24/0.90 34.74/0.91 34.82/0.91 35.00/0.91
Image8 35.94/0.94 34.82/0.93 34.71/0.93 34.66/0.93 35.14/0.94 35.24/ 0.94 35.93/0.94 35.75/0.94 35.92/0.94
Image9 33.25/0.81 32.81/0.79 32.60/0.79 32.32/0.75 33.75/0.80 34.14/0.88 34.24/0.90 34.92/0.89 35.01/0.90
Image10 32.94/0.78 33.52/0.80 32.33/0.83 32.85/0.82 32.84/0.89 32.24/ 0.80 33.96/0.84 35.45/0.82 34.08/0.84
$\tau=30$ Image1 30.82/0.85 30.76/0.84 30.41/0.85 30.36/0.85 30.66/0.86 31.00/0.86 31.30/0.87 31.30/0.87 31.40/0.88
Image2 27.76/0.74 27.01/0.73 26.73/0.73 26.99/0.74 27.30/0.73 27.40/0.73 27.67/0.74 27.77/0.74 28.00/0.74
Image3 31.02/0.81 30.82/0.80 30.74/0.80 30.55/0.80 30.93/0.81 31.10/0.81 31.15/0.81 31.21/0.81 31.22/0.81
Image4 29.45/0.80 29.22/0.79 29.10/0.81 29.00/0.80 29.37/0.81 29.20/0.81 29.57/0.81 29.74/0.81 29.81/0.81
Image5 32.21/0.85 31.02/0.85 31.08/0.85 30.65/0.84 31.34/0.85 31.30/0.85 31.50/0.86 31.32/0.85 31.34/0.85
Image6 31.55/0.93 30.87/0.92 31.01/0.92 30.37/0.93 31.56/0.94 31.61/0.94 31.68/0.94 31.72/0.94 31.50/0.94
Image7 29.23/0.80 28.81/0.79 28.71/0.79 28.94/0.79 29.09/0.80 29.21/0.80 29.32/0.80 29.39/0.80 29.42/0.80
Image8 29.92/0.82 29.56/0.82 29.42/0.81 28.68/0.82 29.82/0.82 29.90/0.82 29.80/0.82 30.06/0.82 30.12/0.83
Image9 29.25/0.80 29.01/0.79 28.80/0.74 28.12/0.69 29.05/0.80 29.24/0.80 29.74/0.81 29.82/0.81 30.02/0.82
Image10 30.94/0.94 30.82/0.93 30.71/0.93 30.66/0.93 31.14/0.94 31.34/ 0.94 31.93/0.94 31.75/0.94 31.92/0.94
$\tau=50$ Image1 28.32/0.82 27.92/0.81 28.01/0.81 28.15/0.82 28.31/0.82 28.30/0.82 28.60/0.82 28.50/0.82 28.60/0.82
Image2 25.23/0.67 25.02/0.66 25.03/0.66 24.94/0.65 25.22/0.67 25.10/0.67 25.47/0.67 25.77/0.67 25.60/0.67
Image3 28.82/0.28 28.75/0.77 28.74/0.76 28.24/0.75 28.85/0.77 28.90/0.77 28.95/0.78 28.91/0.78 28.99/0.78
Image4 27.12/0.74 27.08/0.74 26.60/0.74 26.87/0.74 27.06/0.74 27.20/0.74 27.27/0.75 27.34/0.75 27.51/0.75
Image5 28.56/0.76 28.34/0.76 28.08/0.76 28.01/0.76 28.41/0.77 28.30/0.77 28.50/0.77 28.62/0.78 28.72/0.78
Image6 28.77/0.90 28.53/0.90 28.01/0.89 28.11/0.90 28.52/0.92 28.61/0.92 28.98/0.93 29.02/0.93 29.00/0.93
Image7 26.53/0.75 26.43/0.75 26.21/0.74 26.11/0.74 26.74/0.75 26.71/0.74 26.82/0.75 26.89/0.75 26.92/0.75
Image8 28.12/0.72 27.90/0.71 27.42/0.71 27.62/0.72 28.02/0.72 28.10/0.73 28.10/0.73 28.26/0.73 28.32/0.73
Image9 26.65/0.66 26.71/0.66 26.80/0.68 26.02/0.63 27.05/0.69 27.24/0.70 27.34/0.70 27.62/0.71 28.70/0.74
Image10 27.34/0.70 27.02/0.68 26.81/0.64 26.67/0.62 27.34/0.65 27.64/ 0.70 27.73/0.69 27.75/0.70 28.02/0.72
@@
(c) CBM3D
(d) SRRC
(e) WNNM
(f) DWT

### v2

@@ 5.3.1 Color image denoising without NSS
Image Algorithm pQSTV[37] QWSNM[48] MCNNFNM[33] SRRC[49] QWNNM[15] DWT[7] LRQA-1[12] LRQA-2[12] SLRQA-1 SLRQA-2 SLRQA-3
$\tau=10$ Image1 30.32/0.71 29.00/0.79 30.09/0.79 29.23/0.70 28.06/0.73 27.67/0.61 28.48/0.61 29.21/0.65 31.55/0.76 31.45/0.78 32.86/0.81
Image2 29.21/0.80 30.10/0.79 28.10/0.75 28.56/0.73 28.01/0.70 27.98/0.71 28.44/0.72 28.42/0.72 30.36/0.80 30.19/0.80 31.20/0.82
Image3 30.20/0.78 29.02/0.75 30.12/0.77 28.03/0.60 27.81/0.61 28.21/0.61 28.49/0.64 28.56/0.65 31.33/0.77 31.79/0.78 32.03/0.84
Image4 28.56/0.70 29.20/0.72 28.90/0.69 28.12/0.60 27.90/0.65 28.29/0.66 28.53/0.64 28.55/0.68 30.09/0.74 31.45/0.75 30.36/0.75
Image5 30.92/0.80 30.08/0.78 30.81/0.80 28.82/0.66 28.72/0.64 28.67/0.64 28.85/0.65 28.71/0.67 32.13/0.85 31.18/0.80 32.50/0.87
Image6 28.09/0.70 30.01/0.76 29.40/0.71 28.32/0.60 27.89/0.56 28.02/0.61 28.72/0.65 28.66/0.65 30.75/0.76 30.92/0.80 30.77/0.76
Image7 30.08/0.75 29.60/0.74 30.23/0.75 28.45/0.62 27.32/0.58 28.02/0.59 28.65/0.60 29.04/0.94 30.95/0.73 30.82/0.75 31.10/0.80
Image8 29.20/0.70 30.91/0.71 30.70/0.75 28.01/0.60 28.11/0.59 28.46/0.58 28.56/0.63 28.24/ 0.63 31.23/0.76 30.75/0.74 31.92/0.76
Image9 30.00/0.75 29.80/0.71 29.32/0.72 28.02/0.74 27.98/0.73 27.62/0.71 28.82/0.76 29.10/0.76 30.14/0.70 29.66/0.78 30.80/0.79
Image10 30.86/0.75 30.90/0.75 31.02/0.76 30.52/0.74 29.72/0.72 29.62/0.70 30.90/0.76 31.10/0.79 31.14/0.78 31.26/0.80 31.32/0.81
$\tau=30$ Image1 23.98/0.56 24.80/0.55 25.00/0.57 20.32/0.30 20.41/0.25 20.36/0.28 20.49/0.27 20.80/0.30 25.47/0.66 24.30/0.61 26.40/0.67
Image2 22.80/0.55 23.01/0.57 21.32/0.55 20.01/0.32 20.73/0.32 20.09/0.33 20.44/0.37 20.40/0.37 23.90/0.61 22.77/0.64 24.02/0.64
Image3 22.90/0.54 23.43/0.58 22.03/0.52 20.12/0.21 20.24/0.22 20.55/0.25 20.50/0.27 21.10/0.31 26.15/0.65 25.21/0.64 26.22/0.68
Image4 22.39/0.43 23.09/0.55 23.10/0.55 20.22/0.29 20.10/0.26 20.01/0.29 20.68/0.29 20.20/0.31 23.81/0.59 22.74/0.57 24.21/0.64
Image5 23.03/0.59 24.43/0.60 24.60/0.60 21.02/0.32 20.58/0.33 20.65/0.37 21.23/0.30 21.30/0.33 26.60/0.68 25.32/0.58 26.34/0.68
Image6 21.30/0.59 21.01/0.55 22.30/0.60 20.87/0.33 21.01/0.37 20.37/0.37 20.76/0.36 20.61/0.35 23.95/0.66 21.72/0.64 22.50/0.63
Image7 23.09/0.56 24.33/0.60 24.90/0.61 20.33/0.29 20.71/0.29 20.94/0.22 20.93/0.26 20.21/0.21 25.12/0.63 24.39/0.68 26.42/0.68
Image8 23.33/0.57 24.01/0.58 24.40/0.61 20.56/0.22 20.42/0.24 21.08/0.28 20.77/0.29 20.90/0.29 25.14/0.64 24.06/0.62 26.12/0.67
Image9 24.90/0.60 24.23/0.59 24.89/0.61 24.82/0.58 24.22/0.58 23.62/0.55 24.90/0.58 25.10/0.61 25.30/0.64 25.45/0.68 25.80/0.68
Image10 23.90/0.60 23.84/0.55 24.01/0.62 22.32/0.60 23.82/0.62 23.62/0.61 24.34/0.66 24.90/0.68 24.97/0.66 25.02/0.70 25.68/0.70
$\tau=50$ Image1 18.02/0.32 19.03/0.45 20.30/0.48 17.92/0.17 17.01/0.14 18.01/0.18 17.66/0.17 17.30/0.18 23.33/0.50 23.50/0.51 23.90/0.51
Image2 18.30/0.40 19.02/0.45 18.46/0.42 16.02/0.21 17.03/0.20 17.94/0.25 17.53/0.25 17.10/0.26 22.04/0.48 22.77/0.52 22.06/0.45
Image3 20.01/0.49 20.30/0.52 21.33/0.50 17.75/0.15 17.40/0.21 17.24/0.19 17.56/0.17 17.90/0.18 24.21/0.51 24.91/0.58 24.99/0.60
Image4 20.30/0.39 21.02/0.42 20.80/0.40 17.08/0.14 16.60/0.16 16.87/0.17 17.82/0.19 17.20/0.14 22.23/0.46 22.34/0.47 23.51/0.49
Image5 20.20/0.45 21.03/0.50 20.39/0.51 18.04/0.17 18.08/0.17 18.31/0.18 18.39/0.19 18.30/0.17 21.84/0.52 21.62/0.67 22.72/0.65
Image6 19.32/0.48 19.80/0.50 20.20/0.48 17.53/0.25 18.01/0.26 17.11/0.22 17.75/0.26 18.61/0.29 21.24/0.50 21.02/0.51 22.00/0.53
Image7 20.90/0.49 21.02/0.50 21.30/0.50 17.60/0.17 18.21/0.18 17.11/0.13 18.08/0.17 18.71/0.17 23.57/0.48 23.09/0.45 23.92/0.53
Image8 21.93/0.50 22.03/0.49 22.30/0.51 17.02/0.14 17.42/0.18 17.62/0.19 17.90/0.18 18.10/0.29 23.14/0.48 23.26/0.52 24.32/0.56
Image9 22.30/0.50 21.79/0.51 22.80/0.52 22.02/0.51 21..92/0.49 21.62/0.46 22.34/0.52 22.67/0.55 22.78/0.56 23.02/0.62 23.04/0.63
Image10 22.01/0.50 21.90/0.48 22.34/0.55 21.82/0.50 21.42/0.48 21.32/0.45 22.32/0.51 22.10/0.51 22.75/0.54 22.59/0.52 23.11/0.58
@@
(c) pQSTV
(d) QWSNM
(e) MCNNFNM
(f) SRRC
(g) QWNNM
(h) DWT
(i) LRQA-1
(j) LRQA-2
(k) SLRQA-1
(l) SLRQA-2
(m) SLRQA-3
Figure 4: Color image denoising results of different methods on ”Image5” with $\tau=10$.
@@ 5.3.2 Color image denoising with NSS
For color image denoising problems with NSS, in addition to the methods compared in Section 5.3.1, we also compare SLRQA with CBM3D [44] [^4] The implementation is available from https://webpages.tuni.fi/foi/GCF-BM3D/ released 30 June 2024.. Instead of constructing a denoising model, CBM3D employs block matching and sparse 3D collaborative filtering techniques to denoise the image, a process that cannot be achieved without using NSS. Hence we do not conduct a comparison with the CBM3D algorithm in the experiments of Section 5.3.1.
@@
The PSNR and SSIM values of each image between the competing methods under different noise levels are reported in Table 5. Most proposed SLRQA methods achieve higher PSNR and SSIM values than other methods. Moreover, SLRQA-3 achieves the best results in most cases. The visual comparisons of Image5 with $\tau=30$ are shown in Figure 6 as a typical example. It is shown in Figure 6 that the images recovered by CBM3D and QWNNM methods are still noisy. In addition, the images recovered by DWT, LRQA-1, and LRQA-2 are overly smooth, resulting in the loss of the features of the original images. In contrast, the image recovered by SLRQA-3 can not only remove the noise but also keep the features of the original images.
@@
Table 5: PSNR/SSIM results on color image denoising problem of different methods with NSS. (To highlight the gap between methods, we keep three decimal digits for SSIM.)
@@
Image Algorithm pQSTV [37] QWSNM [48] MCNNFNM [33] CBM3D [44] SRRC [49] QWNNM [20] DWT [7] LRQA-1 [12] LRQA-2 [12] SLRQA-1 SLRQA-2 SLRQA-3
$\tau=10$ Image1 34.78/0.917 35.48/0.921 35.10/0.913 35.71/0.920 35.67/0.919 34.16/0.895 34.62/0.913 35.19/0.919 35.34/0.922 35.68/0.924 35.58/0.921 35.86/0.926
Image2 30.84/0.862 31.98/0.844 31.25/0.839 32.22/0.855 32.02/0.849 31.37/0.862 30.89/0.850 31.94/0.868 31.88/0.866 32.13/0.865 32.19/0.869 32.40/0.872
Image3 36.85/0.934 38.50/0.975 37.89/0.960 37.67/0.948 37.72/0.943 36.37/0.940 36.22/0.934 37.51/0.955 38.62/0.972 38.71/0.973 38.79/0.978 38.83/0.984
Image4 39.79/0.969 40.10/0.975 39.51/0.956 39.36/0.952 39.12/0.948 38.24/0.936 38.89/0.946 39.67/0.955 40.87/0.972 40.99/0.974 41.05/0.976 40.36/0.976
Image5 36.49/0.920 37.05/0.951 36.29/0.943 37.08/0.951 36.82/0.944 35.00/0.921 35.87/0.942 36.48/0.947 36.71/0.945 37.01/0.949 37.18/0.951 37.50/0.954
Image6 35.56/0.924 37.12/0.962 37.08/0.963 37.02/0.969 36.32/0.962 35.42/0.942 36.02/0.959 36.65/0.966 36.66/0.966 37.37/0.969 37.22/0.968 37.27/0.969
Image7 34.11/0.901 35.08/0.916 34.66/0.907 34.25/0.905 34.01/0.894 33.69/0.882 33.92/0.899 34.25/0.904 34.24/0.904 34.68/0.904 34.82/0.905 35.00/0.906
Image8 34.27/0.923 35.77/0.938 35.46/0.934 35.94/0.942 34.82/0.932 34.25/0.910 34.66/0.928 35.14/0.935 35.24/ 0.935 35.84/0.941 35.75/0.940 35.92/0.942
Image9 34.15/0.829 34.83/0.870 33.37/0.826 33.25/0.805 32.81/0.794 32.60/0.788 32.32/0.749 33.75/0.804 34.14/0.884 34.24/0.895 34.92/0.892 35.01/0.902
Image10 33.05/0.812 33.32/0.826 32.80/0.811 32.94/0.780 33.52/0.802 32.84/0.833 32.85/0.818 32.84/0.893 32.24/ 0.804 33.50/0.836 35.45/0.820 34.08/0.842
$\tau=30$ Image1 30.35/0.801 30.21/0.767 29.90/0.817 30.82/0.854 30.76/0.844 30.41/0.854 30.36/0.848 30.66/0.859 31.00/0.862 31.30/0.871 31.30/0.871 31.40/0.877
Image2 27.28/0.722 27.71/0.743 26.67/0.710 27.76/0.740 27.01/0.729 26.73/0.729 26.99/0.737 27.30/0.730 27.40/0.730 27.67/0.740 27.77/0.741 28.00/0.743
Image3 30.05/0.818 30.12/0.802 30.67/0.801 31.02/0.810 30.82/0.801 30.74/0.802 30.55/0.801 30.93/0.811 31.10/0.813 31.15/0.813 31.21/0.814 31.22/0.814
Image4 29.26/0.764 29.52/0.770 28.85/0.767 29.45/0.802 29.22/0.791 29.10/0.806 29.00/0.798 29.37/0.805 29.203/0.81 29.57/0.806 29.74/0.807 29.81/0.809
Image5 30.64/0.846 30.02/0.832 30.35/0.816 32.21/0.847 31.02/0.846 31.08/0.846 30.65/0.837 31.34/0.853 31.30/0.853 31.50/0.859 31.32/0.853 31.34/0.854
Image6 31.35/0.902 30.19/0.823 30.90/0.898 31.55/0.932 30.87/0.923 31.01/0.922 30.37/0.927 31.56/0.941 31.61/0.941 31.68/0.940 31.72/0.941 31.50/0.937
Image7 28.72/0.762 29.02/0.773 28.99/0.778 29.23/0.801 28.81/0.792 28.71/0.790 28.94/0.794 29.09/0.799 29.21/0.801 29.32/0.801 29.39/0.801 29.42/0.802
Image8 30.04/0.808 30.29/0.834 29.36/0.802 29.92/0.820 29.56/0.815 29.42/0.814 28.68/0.817 29.82/0.819 29.90/0.822 29.80/0.818 30.06/0.823 30.12/0.827
Image9 29.00/0.770 29.12/0.782 28.87/0.712 29.25/0.795 29.01/0.794 28.80/0.738 28.12/0.690 29.05/0.799 29.24/0.804 29.74/0.806 29.82/0.805 30.02/0.816
Image10 29.95/0.925 28.71/0.845 29.19/0.713 30.94/0.942 30.82/0.932 30.71/0.929 30.66/0.928 31.14/0.935 31.34/ 0.935 31.93/0.942 31.75/0.940 31.92/0.942
$\tau=50$ Image1 28.07/0.816 28.21/0.767 27.33/0.731 28.32/0.815 27.92/0.810 28.01/0.814 28.15/0.815 28.31/0.818 28.30/0.818 28.60/0.821 28.50/0.819 28.60/0.819
Image2 25.57/0.631 25.36/0.625 24.66/0.630 25.23/0.670 25.02/0.661 25.03/0.660 24.94/0.653 25.22/0.666 25.10/0.666 25.47/0.669 25.77/0.674 25.60/0.670
Image3 28.55/0.728 28.95/0.782 28.57/0.779 28.82/0.778 28.75/0.770 28.74/0.761 28.24/0.754 28.85/0.769 28.90/0.773 28.95/0.781 28.91/0.781 28.99/0.784
Image4 27.01/0.736 27.15/0.712 26.67/0.707 27.12/0.739 27.08/0.740 26.60/0.736 26.87/0.737 27.06/0.742 27.20/0.741 27.27/0.746 27.34/0.746 27.51/0.749
Image5 27.06/0.604 27.82/0.760 27.47/0.697 28.56/0.762 28.34/0.760 28.08/0.756 28.01/0.755 28.41/0.769 28.30/0.767 28.50/0.769 28.62/0.776 28.72/0.775
Image6 28.90/0.907 28.84/0.912 27.81/0.823 28.77/0.900 28.53/0.895 28.01/0.892 28.11/0.899 28.52/0.916 28.61/0.918 28.98/0.930 29.02/0.931 29.00/0.931
Image7 26.04/0.681 26.19/0.692 26.55/0.683 26.53/0.749 26.43/0.745 26.21/0.738 26.11/0.737 26.74/0.747 26.71/0.744 26.82/0.750 26.89/0.751 26.92/0.753
Image8 27.91/0.712 27.74/0.721 26.73/0.702 28.12/0.723 27.90/0.714 27.42/0.714 27.62/0.719 28.02/0.723 28.10/0.729 28.10/0.730 28.26/0.732 28.32/0.732
Image9 27.63/0.664 27.90/0.687 26.19/0.650 26.65/0.655 26.71/0.664 26.80/0.682 26.02/0.629 27.05/0.694 27.24/0.700 27.34/0.701 27.62/0.705 28.70/0.736
Image10 27.07/0.665 27.71/0.702 27.06/0.646 27.34/0.702 27.02/0.682 26.81/0.639 26.67/0.618 27.34/0.652 27.64/ 0.701 27.73/0.687 27.75/0.702 28.02/0.721
@@
(c) pQSTV
(d) QWSNM
(e) MCNNFNM
(f) CBM3D
(g) SRRC
(h) QWNNM
(i) DWT
(j) LRQA-1
(k) LRQA-2
(l) SLRQA-1
(m) SLRQA-2
(n) SLRQA-3
Figure 6: Color image denoising results of tested methods on “Image5” with $\tau=30$.
5.4 Color Image Inpainting
For color image inpainting problems, the compared methods include Bilinear Factorization (BF) methods such as NCALR [25], and Rank Minimization (RM) methods such as SLRI [26], LRQA [12], and QWNNM [20]. 555 The implementations that we use are available from:
QWNNM: https://github.com/Huang-chao-yan/QWNNM
NCARL: https://github.com/hyzhang98/NCARL.
We note that the BF methods require a rank estimate as a prior. For fairness, the SLRI is extended to quaternion representation, which is referred to as QSLRI and performs better than the SLRI method. To fully demonstrate the performance of NCARL, we choose the best PSNR and SSIM from the initial rank range [50, 80, 100, 120, 150].5080100120150[50,\,80,\,100,\,120,\,150].[ 50 , 80 , 100 , 120 , 150 ] . The comparisons with LRQA and QSLRI can be regarded as ablation experiments to present the significance of sparsity and low-rankness prior respectively.
The results of all tested algorithms are shown in Table 6. It is shown therein that SLRQA-NF-3 always enjoys the best performance compared with other methods in terms of PSNR and SSIM. The visual comparisons between SLRQA-NF and all competing inpainting methods on the Image2 with $\chi=0.7$ is shown in Figure 7 as a typical example. Although the NCARL method completes the image, the details of the recovered image are lost, and the recovered images are blurry. The images recovered by LRQA-1 and LRQA-2 methods have a relatively good effect, but the images are still slightly noisy or blurry. In addition, the images recovered by QWNNM are still little noisy. In contrast, SLRQA-NF methods all demonstrate considerable inpainting results and have the highest PSNR and SSIM values.
(a) Image2
(b) Corrupted
(c) NCARL
(d) QWSNM
(e) QWNNM
(f) QSLRI